OpenStudy (anonymous):

A person has 20 coins, all nickels and dimes, worth $1.65. How many nickels are there?

OpenStudy (amistre64):

n + d = p an + bd = q -bn -bd = -bp an +bd = q -------------- n(a-b) = q-bp n = (q-bp)/(a-b)

OpenStudy (anonymous):

Given the sequence 8, 16, 32, 64, ..., which expression would give the thirteenth term?